through athin slit and fils on a A lain beam of completely dark frin observ thai the first pas the Yoe Wim2 at its coma. wide is the slit? of the way that one-quitter what is the intensity of ligh at a point on the screen from the central fringe the fini dak
Answer #1

a. Y=mLR/d

d=0.2*625nm*3/8.24mm

d=0.0455mm

b. d sinA = (0.2*625nm)/(0.0455/4)

sinA= 0.02197

A=0.629

I = Io*cos^2(1.259)

I = 2 w/m^2 * 1

I= 2 w/m^2